OpenStudy (anonymous):

You have a pool that measure 15 yards by 10 yards. You want to build a deck around all four sides of the pool and you want to increase the total area to 250 square yards. What is the width of the deck?

OpenStudy (anonymous):

\[ (15 + x)(10 + x) = 250sq yards \]

OpenStudy (anonymous):

\[x^2 + 25x + 150 = 250 \]

OpenStudy (anonymous):

\[x^2 + 25x - 100 = 0 \]

OpenStudy (anonymous):

Using quadratic formula to solve you get 3.5078 and - 28.5 Therefore the total width of the deck on each side of the length and width of the pool is 3.5078 yards.

OpenStudy (anonymous):

Thank you so much. You helped me a lot.

OpenStudy (anonymous):

No problem,Im glad :D...Tag me or pm me if you need anymore help